# Exploring the Physics of the Plasma Liner Experiment: A Multi-dimensional Study with FLASH, OSIRIS, and HELIOS

### v1

@@ VI Conclusions
FLASH was used to simulate 2D target formation from two colliding magnetized plasma jets, although the actual experiment employs 46 jets. Jet dynamics were briefly discussed, but our primary focus was on the resulting target conditions. With jet propagation conditions within experimental ranges, the hydrogen target plasma reached peak ion temperatures of 40 eV and a volume-averaged electron Hall parameter above unity. The goal of forming a preheated, magnetized, quasi-spherical target was achieved, and we expect results to further improve when more than two target-forming jets are used.
@@
Using both FLASH and OSIRIS in 1D, we studied the effects of initial jet velocity on the liner formation process. The OSIRIS results showed that the jets were most collisional at the lowest velocity (13 km/s) and collisionless at the highest velocity (50 km/s), thus the FLASH results only matched well for low-velocity cases. An intermediate, quasi-collisional regime (the 25 km/s case) may be desirable, as some interpenetration of the jets led to a smooth density profile and may help mitigate hydrodynamic instabilities. We also studied the effects of radiation on liner formation using FLASH (OSIRIS does not include radiation). Results showed that with radiation on, the liner material was much colder and denser (i.e., more collisional), suggesting that FLASH may actually compare well to experimental results of moderate velocities (25 km/s). However, this also suggests that the velocity that results in the desirable quasi-collisional regime may be higher than 25 km/s.
@@
We explored target compression in 1D with FLASH, OSIRIS, and HELIOS. All three codes showed similar implosion dynamics with peak compression near 4.5 s and fuel temperatures around 150 eV. Fuel temperatures in the OSIRIS simulation were overall slightly lower due to kinetic effects (i.e., mixing of the liner and target materials). This suggests that the MHD codes may underestimate the level of mixing depending on the relevant plasma conditions when target compression begins, but it was not an overwhelming effect.
@@
Finally, we conducted high-fidelity 3D target compression simulations with FLASH to study effects of liner perturbations and determine the achievable fuel conditions. The experimental-scale simulation results were qualitatively similar to their 1D counterparts, with stagnation times near 4.0 s and peak ion temperatures in the 150-200 eV range. Using realistically-positioned liner perturbations, we were able to quantify the target’s tolerance to liner non-uniformities. Results point to a 5% liner perturbation level (on density and velocity) as a reasonable upper limit for maintaining a mostly spherical implosion and achieve desirable plasma conditions. An additional, idealized 3D target compression simulation highlighted what may be possible on a larger PLX system. With higher densities, a slightly higher liner velocity, and higher initial target temperature, this reactor-scale simulation reached fusion-relevant conditions with temperatures above 1 keV. All of our 3D target compression cases showed evidence of magnetic field losses due to resistive diffusion, which could potentially be mitigated in future PLX experiments.
@@
Future experimental work on PLX will include further exploration of the parameter space to optimize initial conditions and overall performance. Our simulation results throughout this work have emphasized the key physics at play, providing guidance for that optimization process. We have also helped make the case for a larger PLX system, which our work shows is a theoretically viable path to fusion. There is still plenty of additional simulation work that can be done, as PJMIF is such a rich concept from a physics perspective. Regardless of physics intrigue, from a fusion energy production and engineering perspective, the PLX is an attractive design and one that merits further investigation.
Acknowledgements.
This material is based on work supported by the U.S. Department of Energy (DOE) National Nuclear Security Administration (NNSA) under awards DE-NA0004144 and DE-NA0004147, and under subcontracts no. 630138 and C4574 with Los Alamos National Laboratory. We acknowledge support from the U.S. DOE Advanced Research Projects Agency-Energy (ARPA-E) under Award Number DE-AR0001272. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award FES-ERCAP0028680. The authors acknowledge support from the High-Performance Computing group of the Laboratory for Laser Energetics and the Center for Integrated Research Computing at the University of Rochester. The software used in this work was developed in part by the U.S. DOE NNSA- and U.S. DOE Office of Science-supported Flash Center for Computational Science at the University of Chicago and the University of Rochester. This paper describes objective technical results and analysis. Any subjective views or opinions that might be expressed in the paper do not necessarily represent the views of the U.S. DOE or the United States Government.References
- Ikeda (2007) K. Ikeda, Nuclear Fusion 47, E01 (2007).
- Buttery et al. (2019) R. Buttery, B. Covele, J. Ferron, A. Garofalo, C. Holcomb, T. Leonard, J. Park, T. Petrie, C. Petty, G. Staebler, et al., Journal of Fusion Energy 38, 72 (2019).
- Zylstra et al. (2022) A. Zylstra, O. Hurricane, D. Callahan, A. Kritcher, J. Ralph, H. Robey, J. Ross, C. Young, K. Baker, D. Casey, et al., Nature 601, 542 (2022).
- Abu-Shawareb et al. (2024) H. Abu-Shawareb, R. Acree, P. Adams, J. Adams, B. Addis, R. Aden, P. Adrian, B. Afeyan, M. Aggleton, L. Aghaian, et al., Physical Review Letters 132, 065102 (2024).
- Hsu et al. (2012) S. C. Hsu, T. Awe, S. Brockington, A. Case, J. Cassibry, G. Kagan, S. Messer, M. Stanic, X. Tang, D. Welch, et al., Ieee transactions on plasma science 40, 1287 (2012).
- Thio et al. (2019) Y. F. Thio, S. C. Hsu, F. D. Witherspoon, E. Cruz, A. Case, S. Langendorf, K. Yates, J. Dunn, J. Cassibry, R. Samulyak, et al., Fusion Science and Technology 75, 581 (2019).
- Merritt et al. (2014) E. C. Merritt, A. L. Moser, S. C. Hsu, C. S. Adams, J. P. Dunn, A. Miguel Holgado, and M. A. Gilmore, Physics of Plasmas 21 (2014).
- Langendorf et al. (2018) S. J. Langendorf, K. C. Yates, S. C. Hsu, C. Thoma, and M. Gilmore, Physical Review Letters 121, 185001 (2018).
- LaJoie et al. (2024) A. L. LaJoie, F. Chu, A. E. Brown, S. J. Langendorf, J. P. Dunn, G. A. Wurden, F. Witherspoon, A. Case, M. Luna, J. Cassibry, et al., Physics of Plasmas 31 (2024).
- Yates et al. (2020) K. C. Yates, S. Langendorf, S. C. Hsu, J. Dunn, S. Brockington, A. Case, E. Cruz, F. Witherspoon, Y. Thio, J. Cassibry, et al., Physics of Plasmas 27 (2020).
- LaJoie et al. (2023) A. L. LaJoie, F. Chu, S. Langendorf, J. Cassibry, A. Vyas, and M. Gilmore, Review of Scientific Instruments 94 (2023).
- Ryutov (2009) D. Ryutov, Fusion science and technology 56, 1489 (2009).
- Hsu and Langendorf (2019a) S. C. Hsu and S. J. Langendorf, Journal of Fusion Energy 38, 182 (2019a).
- Chu et al. (2023a) F. Chu, S. J. Langendorf, J. Olson, T. Byvank, D. A. Endrizzi, A. L. LaJoie, K. J. McCollam, and C. B. Forest, Physics of Plasmas 30 (2023a).
- Chu et al. (2023b) F. Chu, A. L. LaJoie, B. D. Keenan, L. Webster, S. J. Langendorf, and M. A. Gilmore, Physical Review Letters 130, 145101 (2023b).
- Langendorf and Hsu (2017) S. J. Langendorf and S. C. Hsu, Physics of Plasmas 24 (2017).
- Hsu and Langendorf (2019b) S. C. Hsu and S. J. Langendorf, Journal of Fusion Energy 38, 182 (2019b).
- Awe et al. (2011) T. J. Awe, C. S. Adams, J. S. Davis, D. S. Hanna, S. C. Hsu, and J. T. Cassibry, Physics of Plasmas 18 (2011).
- Davis et al. (2012) J. Davis, S. Hsu, I. Golovkin, J. MacFarlane, and J. Cassibry, Physics of Plasmas 19 (2012).
- Thoma et al. (2013) C. Thoma, D. Welch, and S. Hsu, Physics of Plasmas 20 (2013).
- Schillo and Cassibry (2020) K. Schillo and J. Cassibry, Physics of Plasmas 27 (2020).
- Thompson and Cassibry (2020) S. Thompson and J. Cassibry, Physics of Plasmas 27 (2020).
- Fryxell et al. (2000) B. Fryxell, K. Olson, P. Ricker, F. X. Timmes, M. Zingale, D. Q. Lamb, P. MacNeice, R. Rosner, J. W. Truran, and H. Tufo, Astrophys. J. Suppl. Ser. 131, 273 (2000).
- Tzeferacos et al. (2015) P. Tzeferacos, M. Fatenejad, N. Flocke, C. Graziani, G. Gregori, D. Q. Lamb, D. Lee, J. Meinecke, A. Scopatz, and K. Weide, High Energy Density Phys. 17, Part A, 24 (2015).
- Lee (2013) D. Lee, J. Computat. Phys. 243, 269 (2013).
- Ji and Held (2013) J.-Y. Ji and E. D. Held, Phys. Plasmas 20, 042114 (2013), https://doi.org/10.1063/1.4801022 .
- Tzeferacos et al. (2017) P. Tzeferacos, A. Rigby, A. Bott, A. R. Bell, R. Bingham, A. Casner, F. Cattaneo, E. M. Churazov, J. Emig, N. Flocke, et al., Phys. Plasmas 24, 041404 (2017).
- Braginskii (1965) S. I. Braginskii, Rev. Plasma Phys. 1, 205 (1965).
- Davies et al. (2021) J. R. Davies, H. Wen, J.-Y. Ji, and E. D. Held, Phys. Plasmas 28, 012305 (2021), https://doi.org/10.1063/5.0023445 .
- Fatenejad et al. (2013) M. Fatenejad, B. Fryxell, J. Wohlbier, E. Myra, D. Lamb, C. Fryer, and C. Graziani, High Energy Density Phys. 9, 63 (2013).
- Orban et al. (2013) C. Orban, M. Fatenejad, S. Chawla, S. C. Wilks, and D. Q. Lamb, arXiv preprint arXiv:1306.1584 (2013).
- Orban et al. (2022) C. Orban, M. Fatenejad, and D. Q. Lamb, Physics of Plasmas 29, 053901 (2022), https://pubs.aip.org/aip/pop/article-pdf/doi/10.1063/5.0079493/16599359/053901_1_online.pdf .
- Sauppe et al. (2023) J. P. Sauppe, Y. Lu, P. Tzeferacos, A. C. Reyes, S. Palaniyappan, K. A. Flippo, S. Li, and J. L. Kline, Physics of Plasmas 30, 062707 (2023), https://pubs.aip.org/aip/pop/article-pdf/doi/10.1063/5.0144517/17992379/062707_1_5.0144517.pdf .
- Meinecke et al. (2014) J. Meinecke, H. W. Doyle, F. Miniati, A. R. Bell, R. Bingham, R. Crowston, R. P. Drake, M. Fatenejad, M. Koenig, Y. Kuramitsu, et al., Nat. Physics 10, 520 (2014).
- Meinecke et al. (2015) J. Meinecke, P. Tzeferacos, A. Bell, R. Bingham, R. Clarke, E. Churazov, R. Crowston, H. Doyle, R. P. Drake, R. Heathcote, et al., Proc. Natl. Acad. Sci. U. S. A. 112, 8211 (2015).
- Li et al. (2016) C. K. Li, P. Tzeferacos, D. Lamb, G. Gregori, P. A. Norreys, M. J. Rosenberg, R. K. Follett, D. H. Froula, M. Koenig, F. H. Seguin, et al., Nat. Commun. 7, 1 (2016).
- Tzeferacos et al. (2018) P. Tzeferacos, A. Rigby, A. F. A. Bott, A. R. Bell, R. Bingham, A. Casner, F. Cattaneo, E. M. Churazov, J. Emig, F. Fiuza, et al., Nat. Commun. 9, 1 (2018).
- Rigby et al. (2018) A. Rigby, F. Cruz, B. Albertazzi, R. Bamford, A. R. Bell, J. E. Cross, F. Fraschetti, P. Graham, Y. Hara, P. M. Kozlowski, et al., Nat. Physics 14, 475 (2018).
- White et al. (2019) T. G. White, M. T. Oliver, P. Mabey, M. Kühn-Kauffeldt, A. F. A. Bott, L. N. K. Döhl, A. R. Bell, R. Bingham, R. Clarke, J. Foster, et al., Nat. Commun. 10, 1 (2019).
- Chen et al. (2020) L. E. Chen, A. F. A. Bott, P. Tzeferacos, A. Rigby, A. Bell, R. Bingham, C. Graziani, J. Katz, M. Koenig, C. K. Li, et al., Astrophys. J. 892, 114 (2020).
- Bott et al. (2021) A. F. A. Bott, P. Tzeferacos, L. Chen, C. A. J. Palmer, A. Rigby, A. R. Bell, R. Bingham, A. Birkel, C. Graziani, D. H. Froula, et al., Proc. Natl. Acad. Sci. U. S. A. 118 (2021).
- Meinecke et al. (2022) J. Meinecke, P. Tzeferacos, J. S. Ross, A. F. A. Bott, S. Feister, H.-S. Park, A. R. Bell, R. Blandford, R. L. Berger, R. Bingham, A. Casner, L. E. Chen, J. Foster, D. H. Froula, C. Goyon, D. Kalantar, M. Koenig, B. Lahmann, C. Li, Y. Lu, C. A. J. Palmer, R. D. Petrasso, H. Poole, B. Remington, B. Reville, A. Reyes, A. Rigby, D. Ryu, G. Swadling, A. Zylstra, F. Miniati, S. Sarkar, A. A. Schekochihin, D. Q. Lamb, and G. Gregori, Science Advances 8, eabj6799 (2022), https://www.science.org/doi/pdf/10.1126/sciadv.abj6799 .
- Fonseca et al. (2002) R. A. Fonseca, L. O. Silva, F. S. Tsung, V. K. Decyk, W. Lu, C. Ren, W. B. Mori, S. Deng, S. Lee, T. Katsouleas, and J. C. Adam, in … Science—ICCS 2002, Vol. 2331, edited by P. M. A. Sloot, A. G. Hoekstra, C. J. K. Tan, and J. J. Dongarra (Springer Berlin Heidelberg, Berlin, Heidelberg, 2002) pp. 342–351.
- Hemker et al. (1999) R. G. Hemker, F. S. Tsung, V. K. Decyk, and W. B. Mori, in Proceedings of the 1999 Particle Accelerator Conference, Vol. 5, edited by A. Luccio and W. MacKay (IEEE, Piscataway, NJ, 1999) pp. 3672–3674.
- Miller et al. (2021) K. G. Miller, R. P. Lee, A. Tableman, A. Helm, R. A. Fonseca, V. K. Decyk, and W. B. Mori, Computer Physics Communications 259, 107633 (2021).
- Nanbu (1997) K. Nanbu, Physical Review E 55, 4642 (1997).
- Takizuka and Abe (1977) T. Takizuka and H. Abe, Journal of Computational Physics 25, 205 (1977).
- Peano et al. (2009) F. Peano, M. Marti, L. O. Silva, and G. Coppa, Physical Review E 79, 025701 (2009).
- Fiuza et al. (2011) F. Fiuza, M. Marti, R. A. Fonseca, L. O. Silva, J. Tonge, J. May, and W. B. Mori, Plasma Physics and Controlled Fusion 53, 074004 (2011).
- MacFarlane et al. (2006) J. MacFarlane, I. Golovkin, and P. Woodruff, Journal of Quantitative Spectroscopy and Radiative Transfer 99, 381 (2006), radiative Properties of Hot Dense Matter.
- Hansen et al. (2015) E. C. Hansen, A. Frank, and P. Hartigan, The Astrophysical Journal 800, 41 (2015).
- Hsu et al. (2015) S. C. Hsu, A. L. Moser, E. C. Merritt, C. S. Adams, J. P. Dunn, S. Brockington, A. Case, M. Gilmore, A. G. Lynn, S. J. Messer, and et al., Journal of Plasma Physics 81, 345810201 (2015).
- Wen et al. (2022) H. Wen, C. Ren, E. C. Hansen, D. Michta, Y. Zhang, S. Langendorf, and P. Tzeferacos, Physics of Plasmas 29, 062706 (2022), https://pubs.aip.org/aip/pop/article-pdf/doi/10.1063/5.0087035/16622426/062706_1_online.pdf .
- Cagas et al. (2023) P. Cagas, J. Juno, A. Hakim, A. LaJoie, F. Chu, S. Langendorf, and B. Srinivasan, Physics of Plasmas 30, 053903 (2023), https://pubs.aip.org/aip/pop/article-pdf/doi/10.1063/5.0135389/17177730/053903_1_5.0135389.pdf .
- Spitzer (1962) L. Spitzer, Physics of Fully Ionized Gases (Interscience Publishers, 1962).
- Shang (1999) J. Shang, Journal of Computational Physics 153, 312 (1999).

### v2

@@ VI Conclusions
FLASH was used to simulate 2D target formation from two colliding magnetized plasma jets, although the actual experiment employs 4-6 jets. Jet dynamics were briefly discussed, but our primary focus was on the resulting target conditions. With jet propagation conditions within experimental ranges, the hydrogen target plasma reached peak ion temperatures of 40 eV and a volume-averaged electron Hall parameter above unity. The goal of forming a preheated, magnetized target was achieved, and we expect results to further improve when more than two target-forming jets are used.
@@
Using both FLASH and OSIRIS in 1D, we studied the effects of initial jet velocity on the liner formation process. The OSIRIS results showed that the jets were most collisional at the lowest velocity ( cm s-1) and collisionless at the highest velocity ( cm s-1), thus the FLASH results only matched the OSIRIS results for low-velocity cases. An intermediate, quasi-collisional regime (the cm s-1 case) may be desirable, as some interpenetration of the jets led to a smooth density profile and may help mitigate hydrodynamic instabilities. We also studied the effects of radiation on liner formation using FLASH, since OSIRIS does not include radiation transport. The simulations showed that radiative losses lead to a liner material that is colder and denser (i.e., more collisional), suggesting that FLASH may actually compare well to experimental results of moderate velocities ( cm s-1). However, this also suggests that the velocity that results in the desirable quasi-collisional regime may be higher than cm s-1.
@@
We explored target compression in 1D with FLASH, OSIRIS, and HELIOS. All three codes showed similar implosion dynamics with peak compression near 4.5 s and fuel temperatures around 150 eV. Fuel temperatures in the OSIRIS simulation were overall slightly lower due to kinetic effects (i.e., mixing of the liner and target materials). This suggests that the MHD codes may underestimate the level of mixing depending on the relevant plasma conditions when target compression begins, but it was not an effect that dramatically changed the outcomes.
@@
Finally, we conducted high-fidelity 3D target compression simulations with FLASH to study the effects of liner perturbations and determined the achievable fuel conditions. The experiment-scale simulation results were qualitatively similar to their 1D counterparts, with stagnation times near 4.0 s and peak ion temperatures in the 150-200 eV range. Using realistically-positioned liner perturbations, we were able to quantify the target’s tolerance to liner non-uniformities. Results point to a 5% liner perturbation level (on density and velocity) as a reasonable upper limit for maintaining a mostly spherical implosion and achieving desirable plasma conditions. An additional, idealized 3D target compression simulation highlighted what may be possible on a larger PLX system: With higher densities, a slightly higher liner velocity, and higher initial target temperature, this reactor-scale simulation reached fusion-relevant conditions with temperatures above 1 keV. All of our 3D target compression cases showed evidence of magnetic field losses due to resistive diffusion, which could potentially be mitigated in future PLX experiments.
@@
Future experimental work on the PLX will include further exploration of the parameter space to optimize initial conditions and overall performance. Our simulation results throughout this work have emphasized the key physics at play, providing guidance for that optimization process. We have also helped make the case for a larger PLX system, which our work shows may be a theoretically viable path to fusion. There is still plenty of additional simulation work that can be done, as the PJMIF is such a physics-rich concept. From a fusion energy production and engineering perspective, the PLX is an attractive design and one that merits further investigation.